Nelson can paint a gym in 7 hours. His assistant, Mike, takes 12 hours to paint the same gym. How long will it take them to paint the gym working together?
Answer by richwmiller(17219)

x/7+x/12=1
12x+7x=84
19x=84
x=84/19=4.42